# Suppression of errors in collectively coded information

### v1

@@ Acknowledgements.
Acknowledgments. We thank Irene Chen, Deborah Fygenson, Paul Higgs, Pepijn Moerman, Erik Winfree, and members of the Murugan and Szostak labs for helpful discussions. This work was supported in part by grants from the Sloan Foundation (19518) and the Gordon and Betty Moore Foundation (11479) to J.W.S. and A.M. J.W.S. is an Investigator of the Howard Hughes Medical Institute. A.M. acknowledges support from NIGMS of the National Institutes of Health under award number R35GM151211. M.J.F. is supported by the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a Schmidt Sciences program. This work was supported by the National Science Foundation through the Center for Living Systems (grant no. 2317138).

### v2

@@ Acknowledgements.
Acknowledgments. We thank Irene Chen, Deborah Fygenson, Paul Higgs, Pepijn Moerman, Erik Winfree, and members of the Murugan and Szostak labs for helpful discussions. This work was supported in part by grants from the Sloan Foundation (19518) and the Gordon and Betty Moore Foundation (11479) to J.W.S. and A.M. J.W.S. is an Investigator of the Howard Hughes Medical Institute. A.M. acknowledges support from NIGMS of the National Institutes of Health under award number R35GM151211. M.J.F. is supported by the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a Schmidt Sciences program. This work was supported by the National Science Foundation through the Center for Living Systems (grant no. 2317138) and by the University of Chicago’s Research Computing Center.
Data Availability. All code to run the simulations is publicly available[54]. Experimental data will be publicly available upon publication.